OpenStudy (anonymous):

Solve system using substitution y-2=x -x=y

OpenStudy (anonymous):

y-2=x -x=y -x-2=x +x +x -2=2x /2 /2 x=-1

OpenStudy (anonymous):

what is y

OpenStudy (anonymous):

was getting to that

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

y-2=-1 +2 +2 y=1. lol its ok

OpenStudy (anonymous):

thanks can you help me with another problem

OpenStudy (anonymous):

-(-1)=1 1=1

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

3y+x=-1 x=-3y

OpenStudy (anonymous):

3y+(-3y)=-1 y=-1 x=-3(-1) x=3 3=-3(-1) 3=3

OpenStudy (anonymous):

i don't want to sound rude but uh can you check your work

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

ok bye